# Deep reinforcement learning-based active flow control of an elliptical cylinder: transitioning from an elliptical cylinder to a circular cylinder and a flat plate

### v2

@@ Abstract
This study investigates the effectiveness of active flow control (AFC) technology supported by deep reinforcement learning (DRL) applied to flows around elliptical cylinders at $Re=100$. We vary the aspect ratio ($Ar$) of the elliptical cylinder from an ellipsoid ($Ar=2.0$) to a circular shape ($Ar=1.0$), and ultimately to a flat plate ($Ar=0$). We utilize the proximal policy optimization (PPO) algorithm to precisely control the mass flow rates of synthetic jets located on both the upper and lower surfaces of the cylinder. The control objective focuses on reducing drag, minimizing lift and suppressing vortex shedding. We examine the robustness and adaptability of DRL-based control techniques across different geometric configurations. Our research findings indicate that, for elliptical cylinders with $Ar$ between 1.75 and 0.75, the reduction in drag coefficient ranges from 0.9% to 15.7%, and the reduction in lift coefficient ranges from 95.2% to 99.7%. Notably, the DRL-based control strategy not only significantly reduces lift and drag, but also completely suppresses vortex shedding while using less than 1% of external excitation energy, demonstrating its efficiency and energy-saving capabilities. Additionally, for $Ar$ from 0.5 to 0, the reduction in drag coefficient ranges from 26.9% to 43.6%, and the reduction in lift coefficient from 50.2% to 68.0%. This reflects the control strategy’s significant reduction in both drag and lift coefficients, while also alleviating vortex shedding. Overall, our findings underscore the adaptability and potential of DRL-based AFC in controlling complex fluid dynamics across diverse geometric configurations.
@@ I INTRODUCTION
Recently, the field of AFC has witnessed revolutionary progress, thanks to breakthroughs in machine learning and deep learning technologies.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020) Deep reinforcement learning (DRL) is an avant-garde technique that combines the principles of deep learning with reinforcement learning to optimize decision-making through interactions with an environment.Arulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2020) Deep learning excels at autonomously extracting intricate features from voluminous datasets, while reinforcement learning concentrates on how to make optimal decisions based on these features by guiding the learning process through rewards.Janiesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015) By utilizing deep neural networks to represent policies or value functions within reinforcement learning, DRL enables machines to tackle complex decision-making challenges and high-dimensional state spaces that were previously considered intractable.Arulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2020) Renowned for its robust non-linear modeling capabilities and adaptive learning mechanisms, DRL has the capability to directly learn complex strategies from raw sensory data without relying on manually engineered features.Dargazany (2021) Furthermore, DRL is inherently self-improving, continuously refining its strategies through ongoing interactions with the environment.Wang et al. (2020) The self-learning and adaptive nature of DRL grants a significant advantage, particularly in resolving decision-making problems in dynamic settings.François-Lavet et al. (2018); Wang et al. (2020) The applications of DRL are vast and encompass various domains, including gaming, autonomous vehicles, robotic control, natural language processing, and computer vision. In the gaming industry, distinguished achievements like AlphaGo’s victory over world champions highlight the potential of DRL.Granter, Beck, and Papke (2017) In the realm of autonomous driving, DRL is employed to optimize route planning and decision-making.Sallab et al. (2017) In the field of robotics, it enables robots to learn complex physical tasks.Zhao, Queralta, and Westerlund (2020) Overall, DRL offers the potential to develop adaptive and intelligent control strategies that can adapt to changing environments and optimize performance in a wide range of applications.
@@
In the academic literature, flow control around blunt bodies using DRL has attracted considerable interest due to its pivotal role in reducing drag, a key aspect of fluid dynamics research. This pursuit carries several important implications, including enhanced energy efficiency, improved performance, noise and vibration reduction, and industrial process optimization. As flow around blunt body tends to involve strong non-linearities such as flow separation and recirculation, traditional control methods often struggle in achieving control objectives whereas DRL models have shown promising results. Pioneered by Rabault et al. Rabault et al. (2019), various researchers such as Tang et al. Tang et al. (2020), Heess et al. Heess et al. (2017), Ren, Rabault, and Tang Ren, Rabault, and Tang (2021), and Wang and Xu Wang and Xu (2024) applied the Proximal Policy Optimization (PPO) algorithm to AFC in the flow around a circular cylinder. Consistently, they achieved an approximate 8% reduction in drag in their validation experiments. Furthermore, Li and ZhangLi and Zhang (2022) ,Wang et al.Wang et al. (2022a), Jia and XuJia and Xu (2024), and Wang et al.Wang et al. (2022b) respectively conducted flow control studies on confined circular cylinders and square cylinders using DRL. In addition, Fan et al.Fan et al. (2020) initiated the application of DRL to AFC experiments, optimizing power efficiency by adjusting the rotational speed of auxiliary circular cylinders positioned behind another circular cylinder. Ren, Wang, and TangRen, Wang, and Tang (2021) proposed an AFC strategy to conceal the hydrodynamic characteristics of circular cylinders, such as strong shear forces and periodic shedding vortices. Studies on flow around constrained circular cylinders under different blockage ratios have also received extensive attention in the field Rabault et al. (2019); Tang et al. (2020); Li and Zhang (2022); Wang et al. (2022a); Wang and Xu (2024). The existing studies have successfully demonstrated the effectiveness of using DRL for flow control, particularly in actively manipulating the flow around circular cylinders and square cylinders.
@@
Although active flow control strategies for both circular and square cylinders are pivotal in various engineering applications for managing aerodynamic and hydrodynamic forcesVignon, Rabault, and Vinuesa (2023); Rabault et al. (2020); Garnier et al. (2021); Razdan and Shah (2022), their extension to more intricate, non-ideal geometries poses significant challenges. The effectiveness of DRL strategies in such contexts remains largely unexplored, marked by uncertainties regarding their capacity to generate physically reasonable and adaptable control measures. Current literature indicates a distinct lack of comprehensive studies on the variability of DRL-based flow control strategies across different geometries, such as from streamlined bodies to various bluff bodies. This research proposes to bridge these gaps by focusing on the flow dynamics around parameterized ellipsoids, examining how variations in the aspect ratio influence flow fields and their corresponding DRL-driven control strategies. By extending the investigation from elongated ellipsoids to circular cylinders and subsequently normal flat plates, this study aims to develop a deeper understanding of the physical validity of DRL approaches and their adaptability across varying body geometries.
This work extends the studies conducted by Wang and XuWang and Xu (2024); Jia and Xu (2024) by applying the DRL-based AFC approach to elliptical cylinders, which present more complex flow characteristics and additional control challenges. The objective is to develop effective control strategies to reduce drag, improve flow stability, and enhance overall performance. The structure of this paper is organized as follows: Sec. II introduces the problem description and provides a comprehensive overview of DRL and its framework for application in AFC. Sec. III presents the main results and analysis. Finally, Sec. IV concludes the paper by summarizing the key findings and discussing their implications.
II PROBLEM DESCRIPTION AND METHODOLOGY
II.1 Numerical simulation
@@
where $\mathbf{u}$, $t$ and $p$ are the velocity vector, time, and pressure, respectively. In this study, the $Re$ is set to 100.
@@
The upstream boundary of the computational domain, designated as $\Gamma_{\text{in}}$, is characterized by a parabolic velocity profile oriented in the $x$-axis, where the boundary conditions are specified such that the transverse velocity component is set to zero, and the velocity along the stream direction shows a parabolic profile distribution. On the downstream boundary, denoted as $\Gamma_{\text{out}}$, a Neumann-type boundary condition is prescribed for the velocity field, ensuring that the stress vector is nullified. On the upper and lower boundaries $\Gamma_{\text{w}}$ , a slip-wall boundary condition is employed, which stipulates that both the velocity component perpendicular to these boundaries and the tangential stress vector component along them are assigned a value of zero. The synthetic jets boundaries are denoted by $\Gamma_{i}(i=1,2)$. The velocity within the jets adheres to a parabolic distribution, with the sign of the jet velocity indicative of fluid suction or injection. The design of the synthetic jets ensures mass flow balance, as expressed by the condition $V_{\Gamma_{1}}=-V_{\Gamma_{2}}$, to maintain mass conservation in the system. Except for the synthetic jets, the non-slip solid wall boundary condition $\Gamma_{\text{w}}$ is adopted for the rest of the cylinder.
In compliance with the above requirements, the system is considered with the following boundary conditions
$-\rho\mathbf{n}\cdot\mathbf{p}+Re^{-1}(\mathbf{n}\cdot\nabla\mathbf{u})$ $=0\quad\text{on }\Gamma_{out},$ (2)
$\mathbf{u}$ $=0\quad\text{on }\Gamma_{w},$
$\mathbf{u}$ $=U\quad\text{on }\Gamma_{in},$
$\mathbf{u}$ $=f_{Q_{i}}\quad\text{on }\Gamma_{i},\quad i=1,2.$
Here, $U$ is the inflow velocity profile. $f_{Q_{i}}$ represents the radial velocity profiles simulating the suction or injection of fluid by the jets. The specific formulas of $U$ and $f_{Q_{i}}$ will be given below.
At the inlet $\Gamma_{\text{in}}$, the inflow velocity $U$ along the $x$-axis is prescribed by a parabolic velocity profile in the form,
$U=U_{m}\frac{(H-2y)(H+2y)}{H^{2}}.$ (3)
Here, $U_{m}$ is the maximum velocity magnitude of the parabolic profile, and $H=4.1D$ represents the total height of the rectangular domain. The mean velocity magnitude is $\overline{U}=2U_{m}/3=1$ and is used to define the Reynolds number $Re=\overline{U}D/\nu$.
The functions $f_{Q_{i}}$ represent radial velocity profiles that simulate the suction or injection mechanisms of the jets. These profiles are selected to ensure a smooth transition to the no-slip boundary conditions on the surfaces $\Gamma_{W}$ of the cylinder. Specifically, $f_{Q_{i}}$ is defined as $A(\theta;Q_{i})(x,y)$, with the modulation relying on the angular coordinate $\theta$ as illustrated in Fig. 2. For the jet characterized by a width of $\omega$ and centered at $\theta_{0}$, which is positioned on the cylinder with radius $R$, the modulation is established as follows:
@@
In this study, the incompressible flow solver is based on the open-source CFD package OpenFOAM, as described by Jasak et al.Jasak (2009); Jasak et al. (2007) OpenFOAM is a widely used and validated CFD software package that provides robust numerical algorithms for solving the Navier-Stokes equations. The solver in OpenFOAM utilizes the finite volume method to discretize the computational domain into a mesh made up of control volumes. To ensure numerical stability, the time step is chosen as $\Delta t$ = 0.0005.
Grid system
Considering an elliptical cylinder computational domain with $Ar=1$, the domain is discretized employing a hybrid meshing approach that integrates both structured (quadrilateral) and unstructured (triangular) meshes. As delineated in Fig. 2, the computational domain is divided into 18,484 mesh cells. Specifically, the area surrounding the elliptical cylinder is discretized using a multi-layer quadrilateral mesh, while the remaining regions are discretized with triangular meshes. Fig.2(a) displays the global discretization scheme for the computational domain, highlighting the comprehensive approach to mesh division. Fig.2(b) provides a detailed view of the mesh partitioning near the elliptical cylinder, where the multi-layered quadrilateral meshes surrounding the cylinder are notably finer compared to the triangular meshes. Fig.2(c) illustrates the division of multi-layered quadrilateral meshes around the elliptical cylinder, emphasizing the precision of discretization near the structure. Finally, fig.2(d) showcases the mesh division around the elliptical cylinder and the integration strategy between quadrilateral and triangular meshes. The method for mesh division for ellipses with $Ar$ values other than 1 follows the same approach as that for $Ar=1$, and is not separately demonstrated.
@@
Quantities of interest
The lift coefficient ($C_{L}$) and drag coefficient ($C_{D}$) are defined as
Here, $F_{L}$ and $F_{D}$ represent the lift and drag forces integrated on the surface of the elliptical cylinder, respectively, and $\rho$ is the fluid density. The Strouhal number $(St)$ is used to describe the characteristic frequency of oscillatory flow phenomena and is defined as follows:
$St=\frac{f_{s}\cdot D}{\overline{U}}$ (6)
Where $f_{s}$ is the shedding frequency calculated based on the periodic evolution of the $C_{L}$. $D$ is the characteristic length of the elliptical cylinder, and $\overline{U}=1$.
@@
Taking the case with an $Ar$ of 1 as an example, this study explores mesh convergence and validates the numerical method utilized. Tab. 1 presents the simulation results using three meshes with differing resolutions, enabling a comparative analysis of the maximum drag coefficient $C_{D,\max}$, the mean of the maximum drag coefficient $C_{D,\text{mean}}$, the maximum lift coefficient $C_{L,\max}$, and the $St$. Furthermore, the results of the numerical simulations are compared with the findings of Rabault et al. Rabault et al. (2019) and Schäfer et al. Schäfer et al. (1996) for validation. The simulations conducted using a coarse mesh revealed significant deviations in the drag and lift coefficients when compared to the results by Rabault et al. Rabault et al. (2019), questioning the accuracy of such numerical simulations. On the other hand, the fine mesh, characterized by a high density of grid points, considerably slows down the training speed for reinforcement learning, presenting a challenge for efficient simulation execution. Our aim is to perform numerical simulations with the least number of mesh cells without compromising on computational accuracy. Consequently, the discretization strategy of the main mesh was adopted for subsequent numerical simulations and reinforcement learning training. This strategy not only satisfies the computational accuracy requirements but also keeps the mesh element count to a manageable level. Comparison of the computational results from the main mesh with those obtained by Rabault et al. Rabault et al. (2019). shows very little difference, indicating a high level of accuracy. Furthermore, the accuracy of the main mesh’s results lies within the upper and lower bounds provided by standard benchmarksSchäfer et al. (1996), confirming that the computational precision of the main mesh is adequate for further training purposes. The above analysis ensures the best balance between numerical accuracy and computational efficiency. The methodology for examining mesh dependency for ellipses with $Ar$ other than 1 is consistent with the procedure established for $Ar=1$.
@@
II.2 Deep Reinforcement learning
Deep learning
Deep learning, a subset of machine learning, employs multi-layer artificial neural networks to model and learn intricate nonlinear relationships. Fig. 3 demonstrates how deep learning captures the representation of input data across multiple layers of a neural network, including the input layer, several hidden layers, and the output layer. Each layer is composed of numerous neural units, with information transmitted between nodes via connection weights. The fundamental concept behind deep learning is its ability to autonomously learn data representations through successive layers. Utilizing the backpropagation algorithm, the network autonomously discerns the transformation process from raw data to sophisticated, high-level abstract features, eliminating the need for manual feature extraction.
Figure 3: DRL combines deep learning’s ability to interpret complex data with reinforcement learning’s strategy to make informed decisions.
Reinforcement learning

### v3

@@ Abstract
This study investigates the effectiveness of active flow control (AFC) technology supported by deep reinforcement learning (DRL) applied to flows around an elliptical cylinder at $Re=100$. We vary the aspect ratio ($Ar$) of the elliptical cylinder from an elliptical cylinder ($Ar=2.0$) to a circular cylinder ($Ar=1.0$), and ultimately to a flat plate ($Ar=0$). We utilize the Proximal Policy Optimization (PPO) algorithm to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of a cylinder to achieve reduction in drag, minimization of lift, and suppression of vortex shedding. The robustness and adaptability of DRL-based control techniques are examined across different geometric configurations. Our research findings indicate that, for elliptical cylinders with $Ar$ between 1.75 and 0.75, the reduction in drag coefficient ranges from 0.9% to 15.7%, and the reduction in lift coefficient ranges from 95.2% to 99.7%. The DRL-based control strategy not only significantly reduces lift and drag, but also completely suppresses vortex shedding while using less than 1% of external excitation energy, demonstrating its efficiency and energy-saving capabilities. Additionally, for $Ar$ from 0.5 to 0, the reduction in drag coefficient ranges from 26.9% to 43.6%, and the reduction in lift coefficient from 50.2% to 68.0%. This reflects the control strategy’s significant reduction in both drag and lift coefficients, while also alleviating vortex shedding. It is expected underscore the adaptability and potential of DRL-based AFC in controlling complex fluid dynamics across diverse geometric configurations.
@@ I INTRODUCTION
Recent advances in machine learning and deep learning technologies have revolutionized the field of artificial fluid control Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020). Deep Reinforcement Learning (DRL) combines deep learning and reinforcement learning to optimize decision-making through interactions with the environment Arulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2020). DRL excels at extracting complex features from large datasets and making optimal decisions based on these features Janiesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015). By using deep neural networks, DRL is capable of handling high-dimensional state spaces and enables machines to tackle previously intractable decision-making challenges Arulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2020). DRL’s nonlinear modeling capabilities and adaptive learning mechanisms allow it to learn complex strategies directly from raw sensory data without manual feature engineering Dargazany (2021). It continuously improves its strategies through interactions with the environment Wang et al. (2020), making it advantageous for dynamic decision-making problems François-Lavet et al. (2018); Wang et al. (2020). DRL has found applications in various fields including gaming, autonomous vehicles, robotics, natural language processing, and computer vision. In particular, it has achieved distinguished success, such as AlphaGo’s victory over world champions in gaming Granter, Beck, and Papke (2017), route planning and decision-making optimization in autonomous driving Sallab et al. (2017), and complex physical tasks learning in robotics Zhao, Queralta, and Westerlund (2020).
@@
In literature, DRL has attracted considerable interest for flow control around blunt bodies, as flow around blunt body tends to involve strong non-linearities such as flow separation and recirculation, traditional control methods often struggle in achieving control objectives whereas DRL models have shown promising results. Pioneered by Rabault et al. Rabault et al. (2019), various researchers such as Tang et al. Tang et al. (2020), Heess et al. Heess et al. (2017), Ren, Rabault, and Tang Ren, Rabault, and Tang (2021), and Wang and Xu Wang and Xu (2024a) applied the Proximal Policy Optimization (PPO) algorithm to AFC in the flow around a circular cylinder. Consistently, they achieved an approximate 8% reduction in drag in their validation experiments. Furthermore, Li and ZhangLi and Zhang (2022) ,Wang et al.Wang et al. (2022a), Wang and XuWang and Xu (2024b) and Wang et al.Wang et al. (2022b)He et al. (2023) respectively conducted flow control studies on confined circular cylinder, square cylinder and airfoil using DRL. In addition, Fan et al.Fan et al. (2020) initiated the application of DRL to AFC experiments, optimizing power efficiency by adjusting the rotational speed of auxiliary circular cylinders positioned behind another circular cylinder. Ren, Wang, and TangRen, Wang, and Tang (2021) proposed an AFC strategy to conceal the hydrodynamic characteristics of circular cylinders, such as strong shear forces and periodic shedding vortices. Studies on flow around constrained circular cylinders under different blockage ratios have also received extensive attention in the field Rabault et al. (2019); Tang et al. (2020); Li and Zhang (2022); Wang et al. (2022a); Wang and Xu (2024a).
@@
Although active flow control strategies for both circular and square cylinders are pivotal in various engineering applications for managing aerodynamic and hydrodynamic forcesVignon, Rabault, and Vinuesa (2023); Rabault et al. (2020); Garnier et al. (2021); Razdan and Shah (2022), their extension to more intricate, non-ideal geometries poses significant challenges. The effectiveness of DRL strategies in such contexts remains largely unexplored, marked by uncertainties regarding their capacity to generate physically reasonable and adaptable control measures. Current literature indicates a distinct lack of comprehensive studies on the variability of DRL-based flow control strategies across different geometries, such as from streamlined bodies to various bluff bodies. Based on the studies conducted by Jia and Xu Wang and Xu (2024a, b), we focus on the development of the DRL technqiue for active flow control around elliptical cylinders, with the objective of developing effective control strategies for drag reduction, flow stability enhancement, and overall performance improvement, addressing the research gap in applying DRL to different geometric shapes of elliptical cylinders. The structure of this paper is organized as follows: Sec. II introduces the problem description and provides a comprehensive procedure for implementing DRL. Sec. III presents the main results and analysis. Finally, Sec. IV concludes the paper by summarizing the key findings and discussing their implications.
II PROBLEM AND METHODOLOGY
II.1 Problem statement
@@
where $\mathbf{u}$, $t$ and $p$ are the velocity vector, time, and pressure, respectively. The Reynolds number, denoted by $Re$, is defined as $Re=\frac{\overline{U}D}{\nu}$, where $\overline{U}=2U_{m}/3=1$ represents the mean velocity magnitude, $U_{m}$ is the maximum velocity, and $\nu$ is the kinematic viscosity. In this study, $Re=100$ is utilized.
@@
The boundary conditions for the computational domain are as follows: The upstream boundary ($\Gamma_{\text{in}}$) is characterized by a parabolic velocity profile, the downstream boundary ($\Gamma_{\text{out}}$) is subjected to a Neumann-type boundary condition to nullify the stress vector, the upper and lower boundaries ($\Gamma_{\text{w}}$) employ a slip-wall boundary condition with zero velocity and tangential stress, and the boundaries of the synthetic jets ($\Gamma_{i}$, where $i=1,2$) are assigned a prescribed parabolic velocity profile representing fluid suction or injection. The synthetic jets feature a parabolic velocity distribution, where the sign of the jet velocity indicates fluid suction or injection. The design ensures mass flow balance through the condition $V_{\Gamma_{1}}=-V_{\Gamma_{2}}$, preserving mass conservation in the system. The non-slip solid wall boundary condition is applied to the rest of the elliptical cylinder except for the synthetic jets.
Mathematically, the boundary conditions are written as:
$\begin{array}[]{cl}-\rho\mathbf{n}\cdot\mathbf{p}+Re^{-1}(\mathbf{n}\cdot\nabla\mathbf{u})=0&\text{on\quad}\;\Gamma_{out},\\ \mathbf{u}=0&\text{on \quad}\Gamma_{w},\\ \;\mathbf{u}=U&\text{on \quad}\Gamma_{in},\\ \;\;\;\mathbf{u}=f_{Q_{i}}&\text{on \quad}\Gamma_{i},\quad i=1,2.\end{array}$ (6)
Here, $\mathbf{n}$ denotes the normal vector, $U$ is the inflow velocity, and $f_{Q_{i}}$ represents the radial velocity profiles simulating the suction or injection of fluid by the jets. At the inlet $\Gamma_{\text{in}}$, the inflow velocity $U$ along the $x$-axis is prescribed by a parabolic velocity profile given by:
$U=U_{m}\frac{(H-2y)(H+2y)}{H^{2}}$ (7)
where $H=4.1D$ represents the total height of the rectangular domain.
In particular, we choose $f_{Q_{i}}=A(\theta;Q_{i})(x,y)$, where the modulation relies on the angular coordinate $\theta$ as illustrated in Fig. 1(b). For the jet, which has a width of $\omega$ and is centered at $\theta_{0}$ on the cylinder with radius $R$, the modulation is defined as follows:
@@
This modulation function ensures a smooth transition to the no-slip boundary conditions on the cylinder surfaces, enabling a comprehensive analysis of the flow characteristics and control challenges associated with the elliptical cylinder geometry.
Quantities of interest
The lift coefficient ($C_{L}$) and drag coefficient ($C_{D}$) are physically important quantities in active flow control, and they are defined as follows:
Here, $F_{L}$ and $F_{D}$ represent the lift and drag forces integrated over the surface of the elliptical cylinder, respectively, and $\rho$ is the fluid density. The Strouhal number ($St$) is used to characterize the characteristic frequency of oscillatory flow phenomena, and it is defined as follows:
$St=\frac{f_{s}\cdot D}{\overline{U}}$ (10)
where $f_{s}$ is the shedding frequency calculated based on the periodic evolution of the $C_{L}$.
II.2 Numerical simulation
@@
We utilize the open-source CFD package OpenFOAM as the incompressible flow solver, following the methodology described by Jasak et al.Jasak (2009); Jasak et al. (2007). This approach is widely employed due to its ability to provide reliable numerical algorithms for solving the Navier-Stokes equations. The solver implemented in OpenFOAM uses the finite volume method to discretize the computational domain into a mesh consisting of control volumes. To ensure numerical stability, a time step size of $\Delta t=0.0005$ is selected in the simulation.
Grid scheme
As an example, we examine a computational domain comprising an elliptical cylinder with an aspect ratio ($Ar=1$) of 1. To discretize this domain accurately, a hybrid meshing approach is adopted, integrating both structured (quadrilateral) and unstructured (triangular) meshes, as shown in Fig. 2. The computational domain is discretized into 18,484 mesh cells, with a multi-layer quadrilateral mesh used for the area surrounding the elliptical cylinder and triangular meshes used for the remaining regions. In detail, Fig. 2(a) provides an overview of the global discretization scheme, demonstrating the comprehensive approach to mesh division employed in the computational domain. Fig. 2(b) offers a closer look at the mesh partitioning near the elliptical cylinder, highlighting the finer multi-layered quadrilateral meshes surrounding the cylinder compared to the triangular meshes. Fig. 2(c) illustrates the division of the multi-layered quadrilateral meshes around the elliptical cylinder, emphasizing the precision in discretization near the structure. Fig. 2(d) illustrates the mesh division strategy around the elliptical cylinder, demonstrating the integration between the quadrilateral and triangular meshes. The method for mesh division for ellipses with $Ar$ values other than 1 follows the same approach as that for $Ar=1$, and is not separately demonstrated.
@@
Similarly, the case where the $Ar$ value is 1 is also used as an example to explore grid convergence and validate the numerical method utilized. Tab. 1 presents simulation results using three different mesh resolutions, allowing for a comparative analysis of the maximum drag coefficient $C_{D,\max}$, the mean of the maximum drag coefficient $C_{D,\text{mean}}$, the maximum lift coefficient $C_{L,\max}$, and the $St$. Furthermore, the results of the numerical simulations are compared with the findings of Rabault et al. Rabault et al. (2019) and Schäfer et al. Schäfer et al. (1996) for validation. The simulations conducted using a coarse mesh reveal significant deviations in the drag and lift coefficients when compared to the results obtained by Rabault et al. Rabault et al. (2019), which raises doubts about the accuracy of such numerical simulations. On the other hand, using a fine mesh with a high density of grid points significantly slows down the training speed for reinforcement learning, posing challenges for efficient simulation execution. It is always desirable to use the minimum number of grid elements for numerical simulation without affecting computational accuracy. As a result, the discretization strategy of the main mesh is adopted for subsequent numerical simulations and reinforcement learning training. The comparison of the computational results obtained from the main mesh with those obtained by Rabault et al. Rabault et al. (2019) reveals minimal differences, indicating a high level of accuracy. Additionally, the accuracy of the main mesh’s results falls within the upper and lower bounds provided by standard benchmarks Schäfer et al. (1996), which confirms that the computational precision of the main mesh is sufficient for further training purposes. The methodology for examining mesh dependency for ellipses with an $Ar$ value other than 1 follows the same procedure established for $Ar=1$.
@@
II.3 Deep reinforcement learning
Model structure